Find the complete integral using Charpit Method : p+q=3pq


1
Expert's answer
2021-01-21T20:08:22-0500

Let's solve this expression:

"f(x, y, z, p,q) = p + q - 3pq = 0" 1)

Charpit's auxilary equations for 1):


"\large\frac{dp}{\large\frac{\delta f}{\delta x} +p\large\frac{\delta f}{\delta z} }=\large\frac{dq}{\large\frac{\delta f}{\delta y} +q\large\frac{\delta f}{\delta z} }=\large\frac{dz}{-p\large\frac{\delta f}{\delta p} -q\large\frac{\delta f}{\delta q} }=\large\frac{dx}{\large\frac{-\delta f}{\delta p} }=\large\frac{dy}{\large\frac{\delta f}{\delta q}}"


"\large\frac{dp}{0+p.0} = \large\frac{dq}{0+q.0} + \large\frac{dz}{-p(1-3q)-q(1-3p)} = \large\frac{dx}{3q-1} = \large\frac{dy}{3p-1}" 2)


Taking the first fraction of (2), we obtain "dp=0"

Integrating it, we get "p = a"

Substituting the value p = a in (1), we get "q = \large\frac{a}{3a-1}"


Now, putting the values of p and q respectively from(3) and (4) in "dz = pdx+qdy," we obtain

"dz= adx +" "\frac{a}{3a-1} dy"

Integrating it, we obtain "z = ax + \large\frac{ay}{3a-1}+b"

Thus, the required comlete integral is "z = ax + \large\frac{ay}{3a-1}+b"